Session: Jun 2005 (session 2)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (C. Bortolotti at Sun Jun 5 06:41:39 2005) Success :

Onsala (Michael Lindqvist at Sun Jun 5 06:46:50 2005) Minor failures :
1. Rather bad weather, rainy from time to time 2. Mark5 hanged, day 154, around 17 UT, lost about 30min 3. According to scan_check (in checkmk5a) there might be problem with the scan (note the "?"): 2005.154.11:18:13.28/scan_check/62,eb030_j0840+1312_no0040,mark4,32,2005y154d9h56m1.185s,?,?,? 4. RFI ==> Bad tsys in a few cases.

Urumqi (yang wen jun at Sun Jun 5 07:26:11 2005) Success :

Torun (K. Borkowski at Mon Jun 6 07:29:57 2005) Minor failures : PCAL
Last 8 hours (since 22:10 on day 155) lost because after crash of one of computers the operator failed to properly set attenuators. Also, there was no phase-cal signal during this experiment.

Westerbork (Tony Foley at Mon Jun 6 11:54:57 2005) Minor failures : WIND
Wb used in phased array mode. AAAAARGH I overwrote the log file by a typo We also had RTC in the array from 06:30 to 09:56 on Friday (The PI requested both RTC and RTD out) On friday evening (16:32) on of the telescopes had a reset caused by nearby lightning strike - so it was out for a few minutes.

Effelsberg (A. Kraus at Tue Jun 7 09:00:35 2005) Minor failures : WIND TSYS SLEW
Thunderstorm: stopped UT1615-1655 (June 3rd). Most Tsys measurements lost due to inadequate slewing time.

Hartebeesthoek (Jonathan Qcuik at Tue Jun 7 09:40:05 2005) Minor failures : RECO RFI
The Mark5A failed to record correctly after an automatic bank switch at d154 23:33 UT though the record pointer did increment - the FS only giving an error at disk_record=off at 00:15 UT. Mark5 manually rebooted and recording restarted at 00:45 UT. Intermittent RFI throughout experiment

Shanghai (Tao An at Tue Jun 7 14:15:21 2005) Success :

Shanghai (Tao An at Tue Jun 7 14:20:47 2005) Minor failures :
we lost some data from UT 07:57:00 to 07:58:00 in scan of No0016. the RF LO is not locked from UT 17:50~17:50 2005.154.23:33:40.28/wx/ 24.2,1007.9,62.0 2005.154.23:39:42.02;"we'v restored wx instrument now. we lost some data from UT 00:14:00 to 00:16:00 of scan no0447. 2005.156.05:06:17.54;"LO unlock at 04h50m, recovered at 05h05m

Jodrell_Bank (Alastair Gunn at Thu Jun 16 10:32:19 2005) Success :
The Lovell Telescope was used. No known problems.

Cambridge (Alastair Gunn at Thu Jun 16 10:32:59 2005) Minor failures :
Several elevation motor trips - odd few minutes lost here and there. No other known problems.